OpenStudy (anonymous):

what are the steps I need to do to simplify 1 over A to the third power?

OpenStudy (fibonaccichick666):

Can you write it out first? (in algebraic terms, the equation button will help you)

OpenStudy (anonymous):

so like 1/a * 1/a* 1/a ?

OpenStudy (fibonaccichick666):

Well, there are a few ways of interpreting this but yes, that is one of the ways

OpenStudy (anonymous):

ok good.

OpenStudy (fibonaccichick666):

so just condense it into 1 fraction

OpenStudy (anonymous):

1/a^3 ?

OpenStudy (fibonaccichick666):

mhmm now the last thing you can do is factor out the cube from the top and bottome

OpenStudy (anonymous):

How do I do that?

OpenStudy (fibonaccichick666):

So \[\frac{1}{a^3}=\frac{1^3}{a^3}\] Do you agree?

OpenStudy (anonymous):

yes.

OpenStudy (fibonaccichick666):

So now if I told you (with a,b,n being any real number): \[\frac{a^n}{b^n}=(\frac{a}{b})^n\] Can you simplify our eq.

OpenStudy (anonymous):

no.

OpenStudy (fibonaccichick666):

Why?

OpenStudy (anonymous):

I dont quite understand it.

OpenStudy (fibonaccichick666):

\[\frac{1}{a^3}=\frac{1^3}{a^3}\] \[\frac{a^n}{b^n}=(\frac{a}{b})^n\] Do they look anything alike?

OpenStudy (anonymous):

Yes except one of them has parentheses.

OpenStudy (fibonaccichick666):

But do you notice what happened to the exponents when the parentheses went around the fraction?

OpenStudy (anonymous):

Oh I see, if you put exponents around it it effects the whole thing

OpenStudy (fibonaccichick666):

yup

OpenStudy (fibonaccichick666):

So what will ours become

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

that was weird

OpenStudy (fibonaccichick666):

o.O?

OpenStudy (anonymous):

I still do not understand how to solve this.

OpenStudy (fibonaccichick666):

No I'm confused haha, you already did it. What are ya missing?

OpenStudy (anonymous):

oh. is that it? (1/a)^3 ?

OpenStudy (fibonaccichick666):

mhmm, I'd list all three though because it could be interpreted all three ways depending on how the original problem was written

OpenStudy (anonymous):

Oh I see thank you very much

OpenStudy (fibonaccichick666):

np